# disposition

### v2

@@ 6.1 Estimators and Asymptotic Distributions
We consider a sequence of binary random variables $(X_{i},Y_{i})$ that are distributed as $(X,Y)=(\mathds{1}_{A},\mathds{1}_{B})$ for all $i\in\mathbb{N}$ and use the shorthand notations $p:=\operatorname{\mathbb{P}}(A)$, $q:=\operatorname{\mathbb{P}}(B)$, $r:=\operatorname{\mathbb{P}}(A\cap B)$ and $\sigma:={Cov}(A,B)=r-pq$. For a given sample of size $n\in\mathbb{N}$, we introduce the natural estimators $\widehat{{p}}_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$, $\widehat{{q}}_{n}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$ and $\widehat{{r}}_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}Y_{i}$ as well as $\widehat{\sigma}_{n}:=\widehat{{r}}_{n}-\widehat{{p}}_{n}\widehat{{q}}_{n}$, $\widehat{m}_{n}^{+}:=\min(\widehat{{p}}_{n},\widehat{{q}}_{n})-\widehat{{p}}_{n}\widehat{{q}}_{n}$ and $\widehat{m}_{n}^{-}:=\widehat{{p}}_{n}\widehat{{q}}_{n}-\max(0,\widehat{{p}}_{n}+\widehat{{q}}_{n}-1)$. Assume that $10<\widehat{{p}}_{n},\widehat{{q}}_{n}<$ 1 holds such that we can define the following plug-in estimators for ${Q}$, ${C}$, and $\phi$ based on Definitions 4.3, 5.1 and 5.3, respectively,
@@
that are well-defined given the condition $10<\widehat{{p}}_{n},\widehat{{q}}_{n}<$ 1, which holds asymptotically with probability one (see (15) for details) by the following assumption.
@@
Let $10<p,q<$ 1 and $\max(0,p+q-1)<r<\min(p,q)$.
Assumption 6.1 is imposed to rule out boundary effects: Violating the conditions $10<p,q<$ 1 would imply that the corresponding estimators $\widehat{{p}}_{n},\widehat{{q}}_{n}$ are almost surely zero or one. The condition $\max(0,p+q-1)=r$ would imply perfect negative dependence and $r=\min(p,q)$ perfect positive dependence such that our estimators for ${Q}$ and ${C}$ (and $\phi$ in the attainable case) would be either plus or minus 1 almost surely, and hence their asymptotic distributions would be degenerate.
@@
Let $(X_{i},Y_{i})_{i\in\mathbb{N}}$ be a stationary ergodic adapted mixingale of size $-1$ and let the “long-run” covariance matrix $\Omega:=\lim_{n\to\infty}\mathrm{Var}\left(n^{-1/2}\sum_{i=1}^{n}\mathbf{W}_{i}\right)$ be strictly positive definite.
In the iid case of Assumption 6.2, $\Omega=\mathrm{Var}\left(\mathbf{W}_{i}\right)$ is positive definite through Assumption 6.1 that assures that the entries in $\mathbf{W}_{i}$ are not perfectly dependent; see the proof of Lemma E.2 for details. Instead, an equivalent condition has to be imposed on the long run variance in the time series case of Assumption 6.3. See White, (2001, Section 5.3) for details on mixingale processes. While time series asymptotics usually come at the cost of more restrictive moment conditions, these are of no concern here due to the boundedness of the binary random variables. We further discuss the time series case in Appendix C.
@@
As ${Q}$ is a continuously differentiable transformation of $p,q$ and $r$ (with Jacobian matrix $J_{g}$ at $(p,q,r)$), the proof of Proposition 6.4 is a straightforward application of the delta method, but is complicated by the need to show non-degeneracies in the Jacobian $J_{g}$.
@@
[…]
@@
    • if $p=1-q$, then $\sqrt{n}\big{(}\widehat{{C}}_{n}-{C}\big{)}\stackrel{{\scriptstyle d}}{{\to}}\Lambda^{-}\begin{pmatrix}V_{4}\\ V_{3}-\mathds{1}\{V_{1}>-V_{2}\}\big{[}V_{1}+V_{2}\big{]}\end{pmatrix}$.
Proposition 6.5 shows that the asymptotic behavior of $\widehat{{C}}_{n}$ is highly non-standard and exhibits discontinuities that depend on the unknown population quantities $\sigma$, $p$ and $q$. The case distinction in whether $\sigma$ (or equivalently ${C}$) is zero, positive or negative arises due to the differing normalizations in the definition of ${C}$: For the case of ${C}=0$, the asymptotic distribution is given by two half-normals with different variances. For $\{{C}>0,\,p\not=q\}$ and $\{{C}<0,\,p\not=1-q\}$, we obtain a standard asymptotic normal distribution. We however get non-standard asymptotic distributions for the cases $\{{C}>0,\,p=q\}$ and $\{{C}<0,\,p=1-q\}$ that arise due to the minimum and maximum operators in the normalizations $m^{+}$ and $m^{-}$.
Figure 3: This figure illustrates the different asymptotic distributions of the estimated Cole coefficient $\widehat{{C}}_{n}$ depending on three choices of the marginal probabilities $(p,q)$ coupled with ${C}\in\{-0.5,0,0.5\}$ in the different subplots. The histograms show the estimated values of $\widehat{{C}}_{n}$ for a sample size of $n=2{,}000$ for $10{,}000$ simulation replications. The plots are augmented with the respective asymptotic densities and the vertical dashed blue lines indicate the true values of ${C}$.
Figure 3 illustrates the five different cases of the asymptotic distribution through simulations. Besides the standard asymptotic normality, we can clearly see the two half-normal distributions for ${C}=0$, where the difference of the variances depends on the specific setting governed by $p$ and $q$. The complicated cases $\{{C}>0,\,p=q\}$ and $\{{C}<0,\,p=1-q\}$ arise in the upper right and middle left subplots. There we also find a bias in the estimates, which is reflected in the asymptotic distributions through the terms $\mathds{1}\{V_{1}>V_{2}\}\big{[}V_{1}-V_{2}\big{]}$ and $\mathds{1}\{V_{1}>-V_{2}\}\big{[}V_{1}+V_{2}\big{]}$, respectively. For ${C}>0$, this bias can be explained by the normalization $\widehat{m}_{n}^{+}=\min(\widehat{{p}}_{n},\widehat{{q}}_{n})-\widehat{{p}}_{n}\widehat{{q}}_{n}>0$, which is negatively biased through $\min(\widehat{{p}}_{n},\widehat{{q}}_{n})$, which tends to underestimate its population counterpart $\min(p,q)$ in the case $p=q$. The argument is similar for the case ${C}<0$.
@@ 6.2 Feasible Inference, Tests and Confidence Intervals
Feasible inference requires consistent estimates of the asymptotic distributions. Denote by $\widehat{J}_{n,g}$, $\widehat{J}_{n,h^{+}}$, $\widehat{J}_{n,h^{-}}$, $\widehat{J}_{n,f}$, $\widehat{J}_{n,l}$, $\widehat{\Lambda}_{n}^{+}$, $\widehat{\Lambda}_{n}^{-}$, $\widehat{m}_{n}^{-}$, $\widehat{m}_{n}^{+}$, and $\widehat{\Delta}_{n}$ the estimated counterparts of the vectors and matrices defined in Propositions 6.4, 6.5 and 6.6 that arise by simply replacing the true probabilities $p$, $q$ and $r$ by their estimators $\widehat{{p}}_{n}$, $\widehat{{q}}_{n}$ and $\widehat{{r}}_{n}$. In the iid case, consistent estimation of $\Omega$ through a plug-in estimator using $\widehat{{p}}_{n},\widehat{{q}}_{n}$ and $\widehat{{r}}_{n}$ in (20) in Appendix E is straight-forward. In the time series case, we employ the well-known “HAC”-estimator of Newey and West, (1987) whose consistency is shown in Proposition C.1. The continuous mapping theorem implies that the empirical “variance” matrices in Propositions 6.4, 6.5 and 6.6 are consistent estimators of their population counterparts, which ensures that the asymptotic distributions are unchanged when the population matrices are replaced with these estimators.
Feasible inference in the form of testing and confidence intervals for ${Q}$ and $\phi$ can be carried out by standard methods based on the above asymptotic normality statements. In contrast, the case distinctions in the asymptotic distribution of $\widehat{{C}}_{n}$ in Proposition 6.5 complicate the construction of tests and confidence intervals as the population quantities ${C}$, $p$ and $q$ are unknown.
@@
Tests for $\mathcal{H}_{0}=\{{C}=0\}$ (or for $\{{C}\leq 0\}$ or $\{{C}\geq 0$}) are straightforward using the respective limiting distribution for ${C}=0$ in Proposition 6.5.222With the same reasons as given after “second” in the following paragraph, we write $\mathcal{H}_{0}=\mathcal{H}_{0}\cap\{\sigma=0\}$ and include a two-sided test for $\{\sigma=0\}$ and obtain the final $p$-value with a Bonferroni correction. When testing for $\mathcal{H}_{c}$ for any $c>0$, we are however uncertain whether we are in the case $\{p=q\}$ or $\{p\not=q\}$ such that a valid testing approach must accommodate both possibilities. Hence, we write $\mathcal{H}_{c}=\mathcal{H}_{=}\cup H_{\not=}$ with ==$\mathcal{H}_{=}=\{{C}=c,\,p=q\}$ and =$\mathcal{H}_{\not=}=\{{C}=c,\,p\not=q\}$, where $\mathcal{H}_{=}$ and $\mathcal{H}_{\not=}$ can be tested using the asymptotically valid $p$-values $p_{=}$ and $p_{\not=}$ that arise from the respective asymptotic distributions in Proposition 6.5. Thus, testing for $\mathcal{H}_{c}$ through the union of these two hypotheses becomes an intersection-union (IU) test, for which $\max(p_{=},p_{\not=})$ is an asymptotically valid $p$-value (Berger,, 1982).
@@
This testing procedure is however overly conservative for two reasons: First, given the predominant case $\{p\not=q\}$, the $p$-value $p_{=}$ is often non-informative and dominates the IU-construction $\max(p_{=},p_{\not=})$. We resolve this by writing $\mathcal{H}_{=}=\mathcal{H}_{=}\cap\mathcal{H}_{p,q}$ with $\mathcal{H}_{p,q}=\{p=q\}$, which we test for by the joint CLT for $(\widehat{{p}}_{n},\widehat{{q}}_{n})$ in Lemma E.2 in the Appendix with $p$-value $p_{p,q}$. Second, the estimated asymptotic variances often differ substantially for positive and negative values of $c$ (see e.g., the middle column of Figure 3), resulting in cases where for a large positive value of ${C}$, we reject many positive values $0<c<{C}$, but cannot reject values $c\leq 0$ due to a higher estimated variance in the negative case. We solve this by writing $\mathcal{H}_{c}=\mathcal{H}_{c}\cap{\mathcal{H}}_{\sigma}$ with ${\mathcal{H}}_{\sigma}=\{\sigma\geq 0\}$, which is a necessary condition for $c>0$. A test for ${\mathcal{H}}_{\sigma}$ with $p$-value $p_{\sigma}$ is easily constructed based on the CLT for $\widehat{\sigma}_{n}$ provided in Lemma E.3 in the Appendix. Both these refinements are union-intersection (UI) tests, for which Bonferroni corrections are necessary. Hence, our final $p$-value is given by $2\min\{\max[2\min(p_{=},p_{p,q}),p_{\not=}],p_{\sigma}\}$ corresponding to the hypothesis $\mathcal{H}_{c}=\{[(\mathcal{H}_{=}\cap H_{p,q})\cup\mathcal{H}_{\not=}]\cap\mathcal{H}_{\sigma}\}$.
For any $c<0$, we use an equivalent construction with ==$\mathcal{H}_{=}=\{{C}=c,\,p=1-q\}$ and =$\mathcal{H}_{\not=}=\{{C}=c,\,p\not=1-q\}$, $\mathcal{H}_{p,q}=\{p=1-q\}$ and $\mathcal{H}_{\sigma}=\{\sigma\leq 0\}$.
@@
In Appendix D, we illustrate the good finite sample properties for the standard confidence intervals for ${Q}$ and $\phi$ and for the more refined construction for ${C}$. For the latter, we also demonstrate the superiority of our approach based on the two auxiliary UI-tests for $\mathcal{H}_{p,q}$ and ${\mathcal{H}}_{\sigma}$ in terms of length of the resulting confidence intervals.

### v3

@@ 6.1 Estimators and Asymptotic Distributions
We consider a sequence of binary random variables $(X_{i},Y_{i})$ that are distributed as $(X,Y)=(\mathds{1}_{A},\mathds{1}_{B})$ for all $i\in\mathbb{N}$ and use the shorthand notations $p:=\operatorname{\mathbb{P}}(A)$, $q:=\operatorname{\mathbb{P}}(B)$, $r:=\operatorname{\mathbb{P}}(A\cap B)$, $\sigma:={Cov}(A,B)=r-pq$, $m^{+}:=\min(p,q)-pq$ and $m^{-}:=pq-\max(0,p+q-1)$. For a given sample of size $n\in\mathbb{N}$, we introduce the natural estimators $\widehat{{p}}_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}$, $\widehat{{q}}_{n}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$ and $\widehat{{r}}_{n}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}Y_{i}$ as well as $\widehat{\sigma}_{n}:=\widehat{{r}}_{n}-\widehat{{p}}_{n}\widehat{{q}}_{n}$, $\widehat{m}_{n}^{+}:=\min(\widehat{{p}}_{n},\widehat{{q}}_{n})-\widehat{{p}}_{n}\widehat{{q}}_{n}$ and $\widehat{m}_{n}^{-}:=\widehat{{p}}_{n}\widehat{{q}}_{n}-\max(0,\widehat{{p}}_{n}+\widehat{{q}}_{n}-1)$. Assume that $0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1$ holds such that we can define the following plug-in estimators for Q{Q}, C{C}, and $\phi$ based on Definitions 4.3, 5.1 and 5.3, respectively,
@@
that are well-defined given the condition $0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1$, which holds asymptotically with probability one (see (15) for details) by the following assumption.
@@
Let $0<p,q<1$ and $\max(0,p+q-1)<r<\min(p,q)$.
Assumption 6.1 is imposed to rule out boundary effects: Violating the conditions $0<p,q<1$ would imply that the corresponding estimators $\widehat{{p}}_{n},\widehat{{q}}_{n}$ are almost surely zero or one. The condition $\max(0,p+q-1)=r$ would imply perfect negative dependence and $r=\min(p,q)$ perfect positive dependence such that our estimators for Q{Q} and C{C} (and $\phi$ in the attainable case) would be either plus or minus 1 almost surely, and hence their asymptotic distributions would be degenerate.
@@
Let $(X_{i},Y_{i})_{i\in\mathbb{N}}$ be a stationary ergodic adapted mixingale of size 1-1 and let the long-run covariance matrix $\Omega:=\lim_{n\to\infty}\mathrm{Var}\left(n^{-1/2}\sum_{i=1}^{n}\mathbf{W}_{i}\right)$ be strictly positive definite.
In the iid case of Assumption 6.2, $\Omega=\mathrm{Var}\left(\mathbf{W}_{i}\right)$ is positive definite through Assumption 6.1 that assures that the components of $\mathbf{W}_{i}$ are not perfectly dependent; see the proof of Lemma E.2 for details. Instead, an equivalent condition has to be imposed on the long run variance in the time series case of Assumption 6.3. See White, (2001, Section 5.3) for details on mixingale processes. While time series asymptotics usually come at the cost of more restrictive moment conditions, these are of no concern here due to the boundedness of the binary random variables. We further discuss the time series case in Appendix C.
@@
As Q{Q} is a continuously differentiable transformation of $p,q$ and $r$ (with Jacobian matrix $J_{g}$ at $(p,q,r)$), the proof of Proposition 6.4 is a straightforward application of the delta method, but is complicated by the need to show non-degeneracies in the Jacobian $J_{g}$.
@@
  • If $\sigma=0$, then $\big(\Delta^{\top}\Omega\Delta\big)^{-1/2}\sqrt{n}\big(\widehat{{C}}_{n}-{C}\big)\stackrel{{\scriptstyle d}}{{\to}}\frac{Z\,\mathds{1}\{Z>0\}}{m^{+}}+\frac{Z\,\mathds{1}\{Z<0\}}{m^{-}}$.
@@
    • if p=1qp=1-q, then $\sqrt{n}\big(\widehat{{C}}_{n}-{C}\big)\stackrel{{\scriptstyle d}}{{\to}}\Lambda^{-}\begin{pmatrix}V_{4}\\ V_{3}-\mathds{1}\{V_{1}>-V_{2}\}\big[V_{1}+V_{2}\big]\end{pmatrix}$.
Proposition 6.5 shows that the asymptotic behavior of $\widehat{{C}}_{n}$ is non-standard and exhibits discontinuities that depend on the unknown population quantities $\sigma$, $p$ and $q$. The case distinction in whether $\sigma$ (or equivalently C{C}) is zero, positive or negative arises due to the differing normalizations in the definition of C{C}: For the case of C=0{C}=0, the asymptotic distribution is given by two half-normals with different variances. For $\{{C}>0,\,p\not=q\}$ and $\{{C}<0,\,p\not=1-q\}$, we obtain a standard asymptotic normal distribution. We however get non-standard asymptotic distributions for the cases $\{{C}>0,\,p=q\}$ and $\{{C}<0,\,p=1-q\}$ that arise due to the minimum and maximum operators in the normalizations $m^{+}$ and $m^{-}$.
Figure 3: This figure illustrates the different asymptotic distributions of the estimated Cole coefficient $\widehat{{C}}_{n}$ depending on three choices of the marginal probabilities $(p,q)$ coupled with ${C}\in\{-0.5,0,0.5\}$ in the different subplots. The histograms show the estimated values of $\widehat{{C}}_{n}$ for a sample size of n=2,000n=2{,}000 for 10,00010{,}000 simulation replications. The plots are augmented with the respective asymptotic densities and the vertical dashed blue lines indicate the true values of C{C}.
Figure 3 illustrates the five different cases of the asymptotic distribution through simulations. Besides the standard asymptotic normality, we can clearly see the two half-normal distributions for C=0{C}=0, where the difference of the variances depends on the specific setting governed by $p$ and $q$. The complicated cases $\{{C}>0,\,p=q\}$ and $\{{C}<0,\,p=1-q\}$ arise in the upper right and middle left subplots. There we also find a bias in the estimates, which is reflected in the asymptotic distributions through the terms $\mathds{1}\{V_{1}>V_{2}\}\big[V_{1}-V_{2}\big]$ and $\mathds{1}\{V_{1}>-V_{2}\}\big[V_{1}+V_{2}\big]$, respectively. For C>0{C}>0, this bias can be explained by the normalization $\widehat{m}_{n}^{+}=\min(\widehat{{p}}_{n},\widehat{{q}}_{n})-\widehat{{p}}_{n}\widehat{{q}}_{n}>0$, which is negatively biased through $\min(\widehat{{p}}_{n},\widehat{{q}}_{n})$, which tends to underestimate its population counterpart $\min(p,q)$ in the case $p=q$. The argument is similar for the case C<0{C}<0.
@@ 6.2 Feasible Inference, Tests and Confidence Intervals
Feasible inference requires consistent estimates of the asymptotic distributions. Denote by $\widehat{J}_{n,g}$, $\widehat{J}_{n,h^{+}}$, $\widehat{J}_{n,h^{-}}$, $\widehat{J}_{n,f}$, $\widehat{J}_{n,l}$, $\widehat{\Lambda}_{n}^{+}$, $\widehat{\Lambda}_{n}^{-}$, $\widehat{m}_{n}^{-}$, $\widehat{m}_{n}^{+}$, and $\widehat{\Delta}_{n}$ the estimated counterparts of the vectors and matrices defined in Propositions 6.4, 6.5 and 6.6 that arise by simply replacing the true probabilities $p$, $q$ and $r$ by their estimators $\widehat{{p}}_{n}$, $\widehat{{q}}_{n}$ and $\widehat{{r}}_{n}$. In the iid case, consistent estimation of $\Omega$ through a plug-in estimator using $\widehat{{p}}_{n},\widehat{{q}}_{n}$ and $\widehat{{r}}_{n}$ in (20) in Appendix E is straight-forward. In the time series case, we employ the well-known “HAC”-estimator of Newey and West, (1987) whose consistency is shown in Proposition C.1. The continuous mapping theorem implies that the empirical covariance matrices in Propositions 6.4, 6.5 and 6.6 are consistent estimators of their population counterparts, which ensures that the asymptotic distributions are unchanged when the population matrices are replaced with these estimators.
Feasible inference in the form of testing and confidence intervals for Q{Q} and $\phi$ can be carried out by standard methods based on the above asymptotic normality statements. In contrast, the case distinctions in the asymptotic distribution of $\widehat{{C}}_{n}$ in Proposition 6.5 complicate the construction of tests and confidence intervals as the population quantities C{C}, $p$ and $q$ are unknown.
@@
Tests for $\mathcal{H}_{0}=\{{C}=0\}$ (or for $\{{C}\leq 0\}$ or $\{{C}\geq 0$}) are straightforward using the respective limiting distribution for C=0{C}=0 in Proposition 6.5.222With the same reasons as given after “second” in the following paragraph, we write $\mathcal{H}_{0}=\mathcal{H}_{0}\cap\{\sigma=0\}$ and include a two-sided test for $\{\sigma=0\}$ and obtain the final $p$-value with a Bonferroni correction. When testing for $\mathcal{H}_{c}$ for any $c>0$, we are however uncertain whether we are in the case $\{p=q\}$ or $\{p\not=q\}$ such that a valid testing approach must accommodate both possibilities. Hence, we write $\mathcal{H}_{c}=\mathcal{H}_{=}\cup H_{\not=}$ with $\mathcal{H}_{=}=\{{C}=c,\,p=q\}$ and $\mathcal{H}_{\not=}=\{{C}=c,\,p\not=q\}$, where $\mathcal{H}_{=}$ and $\mathcal{H}_{\not=}$ can be tested using the asymptotically valid $p$-values $p_{=}$ and $p_{\not=}$ that arise from the respective asymptotic distributions in Proposition 6.5. Thus, testing for $\mathcal{H}_{c}$ through the union of these two hypotheses becomes an intersection-union (IU) test, for which $\max(p_{=},p_{\not=})$ is an asymptotically valid $p$-value (Berger,, 1982).
@@
This testing procedure is however overly conservative for two reasons: First, given the predominant case $\{p\not=q\}$, the $p$-value $p_{=}$ is often non-informative and dominates the IU-construction $\max(p_{=},p_{\not=})$. We resolve this by writing $\mathcal{H}_{=}=\mathcal{H}_{=}\cap\mathcal{H}_{p,q}$ with $\mathcal{H}_{p,q}=\{p=q\}$, which we test for by the joint CLT for $(\widehat{{p}}_{n},\widehat{{q}}_{n})$ in Lemma E.2 in the Appendix with $p$-value $p_{p,q}$. Second, the estimated asymptotic variances often differ substantially for positive and negative values of $c$ (see e.g., the middle column of Figure 3), resulting in cases where for a large positive value of C{C}, we reject many positive values 0<c<C0<c<{C}, but cannot reject values $c\leq 0$ due to a higher estimated variance in the negative case. We solve this by writing $\mathcal{H}_{c}=\mathcal{H}_{c}\cap{\mathcal{H}}_{\sigma}$ with ${\mathcal{H}}_{\sigma}=\{\sigma\geq 0\}$, which is a necessary condition for $c>0$. A test for ${\mathcal{H}}_{\sigma}$ with $p$-value $p_{\sigma}$ is easily constructed based on the CLT for $\widehat{\sigma}_{n}$ provided in Lemma E.3 in the Appendix. Both these refinements are union-intersection (UI) tests, for which Bonferroni corrections are necessary. Hence, our final $p$-value is given by $2\min\{\max[2\min(p_{=},p_{p,q}),p_{\not=}],p_{\sigma}\}$ corresponding to the hypothesis $\mathcal{H}_{c}=\{[(\mathcal{H}_{=}\cap H_{p,q})\cup\mathcal{H}_{\not=}]\cap\mathcal{H}_{\sigma}\}$.
For any $c<0$, we use an equivalent construction with $\mathcal{H}_{=}=\{{C}=c,\,p=1-q\}$ and $\mathcal{H}_{\not=}=\{{C}=c,\,p\not=1-q\}$, $\mathcal{H}_{p,q}=\{p=1-q\}$ and $\mathcal{H}_{\sigma}=\{\sigma\leq 0\}$.
@@
In Appendix D, we illustrate the good finite sample properties for the standard confidence intervals for Q{Q} and $\phi$ and for the more refined construction for C{C}. For the latter, we also demonstrate the superiority of our approach based on the two auxiliary UI-tests for $\mathcal{H}_{p,q}$ and ${\mathcal{H}}_{\sigma}$ in terms of length of the resulting confidence intervals.